some oxygen gas occupies 11.3m3 at 97.2kpa and 76 centigrade. find its volume at 103 kpa and 21 centigrade


1
Expert's answer
2021-01-27T04:40:05-0500

"\frac{P_1V_1}{T_1}" = "\frac{P_2V_2}{T_2}"

P1= 97.2kpa , V1 = 11.3 m3 , T1 = 76⁰C

P2 = 103kpa , V2 = ? , T2 = 21⁰C

"\frac{97.2×11.3}{76}" = "\frac{103×V_2}{21}"


"\frac{1098.36}{76}" = "\frac{103×V_2}{21}"


V2 = 2.94m³